### v3

# OUJ-FTC-21 OCHA-PP-384

Abstract
We apply Nambu non-equilibrium thermodynamics (NNET)—a dynamics with multiple Hamiltonians coupled to entropy-induced dissipation—to paradigmatic far-from-equilibrium systems. Concretely, we construct NNET realizations for the Belousov–Zhabotinsky (BZ) reaction (oscillatory), the Hindmarsh–Rose neuron model (spiking), and the Lorenz and Chen systems (chaotic), and analyze their dynamical and thermodynamic signatures. Across all cases the velocity field cleanly decomposes into a reversible Nambu part and an irreversible entropy-gradient part, anchored by a model-independent quasi-conserved quantity. This construction reproduces cycles, spikes, and strange-attractor behavior and clarifies transitions among steady, periodic, and chaotic regimes via cross-model diagnostics. These results demonstrate that NNET provides a unified, quantitatively consistent framework for oscillatory, spiking, and chaotic non-equilibrium systems, offering a systematic description beyond the scope of linear-response theories such as Onsager’s relations or GENERIC.
1 Introduction
Classical non-equilibrium thermodynamics and self-organization, ranging from de Groot–Mazur[1] and the Glansdorff–Prigogine theory[2] to Nicolis–Prigogine[3] and Haken’s synergetics[4], address irreversibility, stability, and pattern selection. The GENERIC formalism[5, 6, 7, 8] further unifies reversible and irreversible couplings within a bracket structure. In contrast, our NNET makes the reversible–irreversible [^1] In this work, “reversible” does not mean microscopic time-reversal symmetry in the strict sense. Rather, it refers to dynamical reversibility: when the dissipative term is removed, the resulting (dissipationless) flow is invertible so that the admissible backward trajectory is uniquely determined by the forward one. Positivity constraints on variables do not by themselves invalidate this notion; for instance, dissipationless Lotka–Volterra dynamics remains dynamically reversible even under a positive-domain restriction. This is analogous to particle motion described in polar coordinates, which remains reversible despite coordinate or domain constraints. split explicit via Nambu brackets plus an entropy gradient. For rhythms near stable limit cycles, phase reduction originated with Winfree[9] and Kuramoto[10] and has since been systematized[11, 12]. Extensions that incorporate amplitude degrees of freedom via isostables and Koopman theory[13, 14] enlarge the regime beyond weak forcing; for large oscillator ensembles, the Ott–Antonsen ansatz[15] provides low-dimensional closures. Yet these approaches remain essentially local and do not furnish a thermodynamic partition of reversible versus dissipative contributions. Spatially extended pattern formation is grounded in reaction–diffusion theory from Turing[16] to modern reviews[17, 18]. For chemical oscillations, the Belousov-Zhabotinsky (BZ) reaction is canonically treated by the Oregonator mechanism and comprehensive monographs[19, 20, 21]. On the life-science side, spiking/bursting spans biophysically grounded models[22], geometric classifications[23], and modern syntheses[24]. These literatures motivate a unified far-from-equilibrium decomposition, which we supply via NNET.
To address such systems, we have proposed Nambu Non-equilibrium Thermodynamics (NNET), a framework based on Nambu brackets that couples multiple Hamiltonians with entropy-induced dissipation. In this formulation, the velocity field separates cleanly into a reversible Nambu component and an irreversible entropy-gradient component, organized by a model-independent quasi-conserved quantity (i.e., a slowly varying variable that behaves as conserved on the relevant timescale, such as catalyst concentration in chemical oscillations or the bursting variable in neuron models; see [25] for a formal definition).
Building on the original study “Fluctuating Non-linear Non-equilibrium System in Terms of Nambu Thermodynamics”[26], which integrates reversible Nambu dynamics[27] with entropy-driven dissipation, this paper further develops the framework by broadening and strengthening its applications to concrete models. As contributions of this paper, by analyzing the time evolution of representative models—the Belousov–Zhabotinsky reaction[29, 30] (oscillations), the Hindmarsh–Rose neuron[31] (spiking), and the Lorenz[32] and Chen[33] systems (chaos)—within the Nambu formalism, we demonstrate how Hamiltonian-derived conservative dynamics and entropy-driven dissipation coexist and interact. This establishes that such diverse non-equilibrium behaviors can be understood at least in the paradigmatic cases studied here as manifestations of Nambu Non-equilibrium Thermodynamics (NNET), providing a unifying description beyond the scope of linear-response theories such as those of Onsager[34] or GENERIC[6, 7, 8], in particular by accommodating regimes where entropy can locally decrease as part of the reversible–irreversible interplay.
The remainder of this paper is organized as follows: In Section 2, we describe the construction of Hamiltonians and entropy within the Nambu formalism, restricting throughout to the case $N=3$ for concreteness. The general case of arbitrary $N$ has already been discussed in [28]. Section 3 presents practical applications and numerical analyses for the BZ reaction, the Hindmarsh–Rose[31] model, the Lorenz system[32], and the Chen system[33]. Finally, in Section 4, we conclude with a summary and discussion of future directions. Appendices collect supplementary materials: NNET foundations in summary (for the present 3D applications), a minimal toy model reproducing cycle–spike transitions, a near-harmonic approximation for the BZ limit cycle, and spike onset geometry analytical details.
2 Approaches for Constructing the Hamiltonians and Entropy
2.1 General evolution law of NNET
In this section, we provide an overview of the Nambu Non-equilibrium Thermodynamics (NNET) framework. The Nambu Non-equilibrium Thermodynamics (NNET) framework couples two essential flows: a reversible, volume-preserving flow generated by Nambu dynamics with multiple Hamiltonians, and an irreversible flow driven by the gradient of an entropy-like potential [^2] In Onsager’s near-equilibrium theory, this $S$ coincides with the thermodynamic entropy, but in general it can differ.($S$). This dual structure cleanly partitions the dynamics into conservative and dissipative contributions and furnishes a thermodynamic interpretation of the latter. In what follows, we first present the general evolution law (here restricted to ($N=3$), then give constructive procedures for identifying the Hamiltonians and the entropy [^3] This generalized $S$ can also be interpreted as a Lyapunov function or a dissipation potential.($S$) for a given system, and finally illustrate the recipe on concrete models. A detailed axiomatic foundation is provided in [25].
We begin by recalling the basic objects of the NNET formulation and then state the general evolution law. In NNET, a system is described by an $N$-dimensional thermodynamic state vector $x=(x^{1},\dots,x^{N})$. The dynamics is decomposed into (i) a reversible, volume-preserving flow generated by multiple Hamiltonians $H_{1},\dots,H_{N-1}$ via a Nambu bracket, and (ii) an irreversible flow driven by the gradient of an entropy-like potential $S$. Accordingly, the time evolution of the state variables is given by
$\dot{x}^{i}\;=\;-\{H_{1},H_{2},\cdots,H_{N-1},x^{i}\}_{\mathrm{NB}}\;+\;\partial_{i}S,\qquad(i=1,\dots,N),$ (1)
where the Nambu bracket $\{\cdots\}_{\mathrm{NB}}$, originally introduced by Nambu[27] and defined as a Jacobian,
$\{A_{1},A_{2},\cdots,A_{N}\}_{\mathrm{NB}}\equiv\frac{\partial(A_{1},A_{2},\cdots,A_{N})}{\partial(x_{1},x_{2},\cdots,x_{N})},$ (2)
governs the conservative part of the dynamics, while the gradient term represents the dissipative contribution. Here and below, $\partial^{\,i}S$ denotes the $i$-th component of $\nabla S$.
For a general observable $O=O(x)$, the same decomposition yields
$\dot{O}\;=\;-\{H_{1},H_{2},\cdots,H_{N-1},O\}_{\mathrm{NB}}\;+\;(\nabla S\cdot\nabla O).$ (3)
Equations (1)–(3) define an autonomous system of $N$ ordinary differential equations,
$\dot{x}^{i}\;=\;v^{i}(x),\qquad(i=1,\dots,N),$ (4)
with velocity field $v(x)$. Since the Nambu-generated flow is divergence-free, the entropy $S$ is particularly accessible through the identity
$\nabla\cdot v(x)\;=\;\nabla^{2}S(x).$ (5)
It should be emphasized that the “entropy” $S$ in NNET does not necessarily coincide with the thermodynamic entropy in the sense of Onsager’s near-equilibrium theory. Here the word “entropy” is used in a generalized sense, as a potential generating dissipation, rather than strictly the thermodynamic entropy. This generalized entropy inherits the role of ensuring non-negative dissipation, analogous to the second law in near-equilibrium thermodynamics. When one of the Hamiltonians is identified with $S$ and the system is close to equilibrium, the formulation reduces to the Onsager framework. In general, however, $S$ should be regarded as a potential that generates dissipation and can differ from the original entropy of the nonlinear system; see [28] for discussion. A deeper physical characterization of this generalized entropy via a variational principle is an important subject for future work.
Although the formulation is valid for a general $N$, in this paper we confine our analysis to the case $N=3$, as all subsequent applications (such as the BZ reaction and the H–R model) belong to this class. In this setting, we employ two Hamiltonians, $H_{1}$ and $H_{2}$, and represent the state vector as $x=(X,Y,Z)$.
2.2 Constructive procedure for Hamiltonians and entropy
In practice, once the general NNET evolution law has been specified, the next task is to construct explicit forms of the Hamiltonians and the entropy that reproduce a given set of evolution equations. To this end, we begin with a Helmholtz-type decomposition of the velocity field and then identify the Hamiltonian (incompressible) and entropy (gradient) parts in a constructive manner.
The velocity field is decomposed into an incompressible part governed by Nambu dynamics and a compressible part derived from the entropy gradient,
$v(x)\;=\;\nabla\times\psi\;+\;\nabla S,$ (6)
where, in this subsection, $\nabla$ denotes the spatial gradient (so that $\nabla\times$ and $\nabla\cdot$ are the curl and divergence, respectively) and we work in $N=3$ dimensions. The vector potential $\psi$ is chosen as
$\psi\;=\;\tfrac{1}{2}\,\bigl(H_{1}\overleftrightarrow{\nabla}H_{2}\bigr)\;\equiv\;\tfrac{1}{2}\,\bigl(H_{1}\nabla H_{2}-H_{2}\nabla H_{1}\bigr),$ (7)
so that $\nabla\times\psi=\nabla H_{1}\times\nabla H_{2}$ and the incompressible component is generated by the Nambu bracket $\{H_{1},H_{2},\cdot\}_{\mathrm{NB}}$. Equation (6) thus corresponds to the standard Helmholtz decomposition into a volume-preserving part and a gradient part. Moreover, by Darboux’s theorem, the solenoidal component can always be represented in terms of Nambu brackets with appropriately chosen Hamiltonians; see [28] for details. Consistency with the NNET evolution law further requires the compatibility condition $\nabla\!\cdot\!v(x)=\nabla^{2}S(x)$ (cf. Eq. (5)), which we use to determine $S$ from a given $v$.
To simplify the construction in applications, we often adopt
$H_{2}\;=\;Z,$ (8)
where $Z$ is treated not as an exactly conserved quantity but as a pseudo-conserved variable: it remains approximately constant over the relevant timescale. For instance, in chemical reaction systems, the concentration of a catalyst or intermediate species may oscillate periodically and serve as such a variable. Introducing a pseudo-conserved quantity in this way frequently streamlines the analysis of complex non-equilibrium systems.
Within this assumption, $H_{1}$ and $S$ can be obtained directly from the differential equations:
$\begin{cases}\dot{x}^{1}=2\partial_{2}H_{1}+\partial_{1}S,\\ \dot{x}^{2}=-2\partial_{1}H_{1}+\partial_{2}S,\\ \dot{x}^{3}=\partial_{3}S.\end{cases}$ (9)
These lead to the following compatibility conditions between $H_{1}$ and $S$:
$2\partial_{2}^{2}H_{1}+2\partial_{1}^{2}H_{1}$ $=$ $\partial_{2}\dot{x}^{1}-\partial_{1}\dot{x}^{2},$ (10)
$2\partial_{3}\partial_{2}H_{1}$ $=$ $\partial_{3}\dot{x}^{1}-\partial_{1}\dot{x}^{3},$ (11)
$-2\partial_{3}\partial_{1}H_{1}$ $=$ $\partial_{3}\dot{x}^{2}-\partial_{2}\dot{x}^{3}.$ (12)
These relations can also be obtained from the variational structure of the system. The starting point is the constitutive equation:
$\dot{x}^{i}=-\{H_{1},H_{2},\dots,H_{N-1},x^{i}\}_{\mathrm{NB}}+\partial_{i}S.$ (13)
This can be rewritten in differential form as:
$dS=\delta_{ij}\left(v^{i}+\{H_{1},\dots,H_{N-1},x^{i}\}_{\mathrm{NB}}\right)dx^{j}.$ (14)
Taking the exterior derivative of both sides yields the following consistency condition:
$\delta_{ij}\left(\frac{\partial v^{j}}{\partial x^{k}}+\frac{\partial}{\partial x^{k}}\{H_{1},\dots,H_{N-1},x^{j}\}_{\mathrm{NB}}\right)dx^{k}\wedge dx^{i}=0.$ (15)
If the Hamiltonians are selected so as to satisfy this condition, the remaining terms in the constitutive equation can be written as an exact differential, namely, the gradient of an entropy function $S$.
We now turn to a chemical reaction system characterized by three dynamical variables, $X,Y$, and $Z$. Let $Z$ denote the concentration of a catalyst. In many such systems, the catalyst concentration exhibits periodic oscillations, which makes it natural to regard $Z$ as the second Hamiltonian:
$H_{2}=Z.$ (16)
Substituting this into Eq. (15) yields the following condition:
$\left(2\epsilon_{kj3}\frac{\partial^{2}H_{1}}{\partial x^{i}\partial x^{j}}-2\epsilon_{ij3}\frac{\partial^{2}H_{1}}{\partial x^{k}\partial x^{j}}+\partial_{k}\dot{x}_{i}-\partial_{i}\dot{x}_{k}\right)dx^{k}\wedge dx^{i}=0$ (17)
from which the relations in Eq. (10) immediately follow.
As we will demonstrate in subsequent sections, this condition can be explicitly satisfied in concrete models such as the BZ reaction and the Hindmarsh–Rose model.
3 Application
In this section, we demonstrate that non-equilibrium systems far from equilibrium can be described within a unified framework of Nambu non-equilibrium thermodynamics. As illustrative cases, we analyze the BZ reaction, representing temporal oscillations as a prototypical dissipative structure, and the H–R model, which captures spike–burst behavior of the membrane potential from the perspective of NNET. We then extend the analysis to chaotic dynamical systems, examining both the Lorenz and Chen systems within the same framework.
3.1 BZ reaction
As a representative periodic system, we reconstruct the Belousov–Zhabotinsky (BZ) reaction within the NNET framework and discuss the time dependence of the Hamiltonians ($H_{i}$) and the entropy ($S$). The BZ reaction[29, 30] is a typical case of a time-oscillating chemical reaction. The BZ reaction is a prototypical example of a non-equilibrium chemical oscillation, in which the concentrations of intermediates exhibit periodic temporal changes accompanied by striking color oscillations. As emphasized in the pioneering works of Prigogine and coworkers, the BZ reaction has long been regarded as an archetype of dissipative structures that form far from equilibrium, providing key insights into the emergence of temporal order in chemical systems.
The simplest and most widely used reduced mechanism for the BZ reaction is the Oregonator[19]. It idealizes the chemistry as an interplay of (i) autocatalytic production and consumption of the activator ($X$), (ii) inhibition by bromide ($Y$), and (iii) redox cycling of the catalyst ($Z$) driven by the organic substrate ($B$). Concretely, the elementary steps are written as
$A+Y\;\xrightarrow{k_{1}}\;X+P,$ (18)
$X+Y\;\xrightarrow{k_{2}}\;2P,$ (19)
$A+X\;\xrightarrow{k_{3}}\;2X+2Z,$ (20)
$2X\;\xrightarrow{k_{4}}\;A+P,$ (21)
$B+Z\;\xrightarrow{k_{5}}\;h\,Y.$ (22)
Here we follow the standard identification of species:
$A=\mathrm{BrO_{3}^{-}},\quad X=\mathrm{HBrO_{2}},\quad Y=\mathrm{Br^{-}},\quad Z=\mathrm{Ce^{4+}},\quad P=\mathrm{HOBr},\quad B=\mathrm{CH_{2}(COOH)_{2}},$
where ($k_{1},\dots,k_{5}$) are reaction rate constants and ($h$) is an adjustable stoichiometric factor. In typical experimental conditions the concentrations of ($A$), ($B$), and ($P$) are maintained (buffered or in large excess) and are therefore treated as constants.
Then, the equations of the time evolution of the BZ model become the equations of motion for three coordinates or three concentrations ($X$, $Y$, and $Z$), given by
$\frac{dX}{dt}$ $=$ $k_{1}AY-k_{2}XY+k_{3}AX-2k_{4}X^{2},\$(23)
$\frac{dY}{dt}$ $=$ $-k_{1}AY-k_{2}XY+hk_{5}BZ,\$(24)
$\frac{dZ}{dt}$ $=$ $2k_{3}AX-k_{5}BZ.$ (25)
These chemical reactions are far from equilibrium and are not obtained from the Onsager’s variational principle.
Following the constructive procedure outlined in Section 2.2, we now determine the Hamiltonians $H_{i}$ and the entropy $S$ for the BZ reaction. Following the methodology of Section 2.2, consider $Z$ as the catalyst [^4] Here, the catalyst is interpreted as a quasi-conserved quantity., $H_{2}$.
Then, we obtain
$d\{H_{1},Z,x^{i}\}\wedge dx^{i}=d\dot{x}^{i}\wedge dx^{i}.$ (26)
If we insert the concrete form from the development equation for velocity, $H_{1}$ must satisfy the following relation:
$\frac{\partial^{2}H_{1}}{\partial X^{2}}+\frac{\partial^{2}H_{1}}{\partial Y^{2}}$ $=$ $k_{1}A-k_{2}(X-Y),$ (27)
$\frac{\partial^{2}H_{1}}{\partial Z\partial X}$ $=$ $hk_{5}B,$
$\frac{\partial^{2}H_{1}}{\partial Z\partial Y}$ $=$ $2k_{3}A.$
From this,
$\frac{\partial H_{1}}{\partial X}$ $=$ $hk_{5}BZ+f(X,Y),$ (28)
$\frac{\partial H_{1}}{\partial Y}$ $=$ $2k_{3}AZ+g(X,Y),$
$\frac{\partial f(X,Y)}{\partial X}+\frac{\partial g(X,Y)}{\partial X}$ $=$ $k_{1}A-k_{2}(X-Y).$
In this context, $f$ and $g$ are regarded as functions of $X$ and $Y$. By substituting these into the above formula,
$f(X,Y)=\frac{1}{2}k_{2}X^{2}$ (29)
and
$g(X,Y)=-k_{1}AY+\frac{k_{2}}{2}Y^{2}.$ (30)
Therefore, solving for $H_{1}$ yields
$H_{1}=hk_{5}BZX-\frac{1}{2}k_{1}AY^{2}+\frac{k_{2}}{6}Y^{3}+2k_{3}AYZ+\frac{1}{6}k_{2}X^{3}.$ (31)
The time evolution of the non-Hamiltonian part is necessarily in the form of an entropy gradient. To derive the entropy $S$, we use the decomposition of the velocity field $v$ in Eq. (6) into an incompressible Nambu part and a gradient part, namely., $v=v^{(H)}+\nabla S$. Concretely, after determining $H_{1}$, we define the Nambu (Hamiltonian) contribution to the velocities by
$v^{(H)}=\bigl(-\{H_{1},H_{2},X\}_{\mathrm{NB}},\,-\{H_{1},H_{2},Y\}_{\mathrm{NB}},\,-\{H_{1},H_{2},Z\}_{\mathrm{NB}}\bigr),$ (32)
and the residual by $v^{(S)}:=\dot{x}-v^{(H)}$. The consistency relation Eq. (15) ensures that this residual is an exact gradient, $v^{(S)}=\nabla S$, which is also equivalent to $\nabla\cdot v=\nabla^{2}S$ in Eq. (5). Integrating $\partial_{i}S=v^{(S)}_{i}$ (up to an additive constant) yields the explicit expression. This construction follows [25, 28].
$S=$ $-\frac{k_{2}}{2}XY^{2}-\frac{k_{2}}{2}X^{2}Y$ (33)
$-\frac{2}{3}k_{4}X^{3}+\frac{1}{2}k_{3}AX^{2}-\frac{1}{2}k_{1}AY^{2}$
$-\frac{k_{5}}{2}BZ^{2}+2k_{3}AXZ.$
The split follows directly from the decomposition $v=\nabla\times\psi+\nabla S$ in Eq. (6). For any observable $O$, Eq. (3) reads
$\dot{O}=-\{H_{1},H_{2},O\}_{\mathrm{NB}}+(\nabla S\cdot\nabla O),$ (34)
which we denote by $\partial_{t}^{(H)}O$ and $\partial_{t}^{(S)}O$, respectively. The first term is an incompressible Nambu flow that preserves phase-space volume and encodes the conservative, multi-Hamiltonian structure; the second is a compressible gradient flow generated by $S$ that accounts for dissipation and entropy production. This Hamiltonian–entropy split is precisely the far-from-equilibrium extension in [25], and it is consistent with the existence result discussed in [28]. Operationally, we compute $\partial_{t}^{(H)}$ from $\{H_{1},H_{2},\cdot\}$ and set $\partial_{t}^{(S)}$ by the residual $(\nabla S\cdot\nabla)$.
We divide time evolution into Hamiltonian part and entropy part:
$\frac{d}{dt}=\partial_{t}^{(H)}+\partial_{t}^{(S)}.$ (35)
Then, we obtain the time evolution induced by the Hamiltonians, as follows:
$\partial_{t}^{(H)}X$ $=$ $-\{H_{1},H_{2},X\}_{\mathrm{NB}}=k_{1}AY+\frac{k_{2}}{2}Y^{2}-2k_{3}AZ,\$(36)
$\partial_{t}^{(H)}Y$ $=$ $-\{H_{1},H_{2},Y\}_{\mathrm{NB}}=\frac{k_{2}}{2}X^{2}+hk_{5}BZ,\$(37)
$\partial_{t}^{(H)}Z$ $=$ $-\{H_{1},H_{2},Z\}_{\mathrm{NB}}=0,$ (38)
while the time evolution induced by the entropy is given by
$\partial_{t}^{(S)}X$ $=$ $\{S,Y,Z\}_{\mathrm{NB}}=k_{3}AX-2k_{4}X^{2}-k_{2}XY-\frac{k_{2}}{2}Y^{2}+2k_{3}AZ,\$(39)
$\partial_{t}^{(S)}Y$ $=$ $\{S,Z,X\}_{\mathrm{NB}}=-\frac{k_{2}}{2}X^{2}-k_{1}AY-k_{2}XY,\$(40)
$\partial_{t}^{(S)}Z$ $=$ $\{S,X,Y\}_{\mathrm{NB}}=2k_{3}AX-k_{5}BZ.$ (41)
By combining these two types of time evolution, we can obtain the overall time evolution of the BZ reaction. Next, we examine the time evolution of entropy, which can be expressed as follows:
$\dot{S}$ $=$ $\partial_{t}^{(H)}S+\partial_{t}^{(S)}S,\$(42)
$\partial_{t}^{(H)}S$ $=$ $-\{H_{1},H_{2},S\}_{\mathrm{NB}},\$(43)
$\partial_{t}^{(S)}S$ $=$ $+\{S,X,Y\}_{\mathrm{NB}}^{2}+\{S,Y,Z\}_{\mathrm{NB}}^{2}+\{S,Z,X\}_{\mathrm{NB}}^{2}.$ (44)
The entropic part of the entropy time evolution is always positive, reflecting the essence of entropy production. In contrast, the Hamiltonian part may become negative, as it corresponds to reversible dynamics. Such negative contributions can arise, for example, in oscillatory reactions like the BZ reaction or in non-equilibrium steady states, where reversible energy or matter redistribution leads to a temporary decrease in entropy.
Now
$\{H_{1},H_{2},S\}_{\mathrm{NB}}=\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{t}^{(H)}X\partial_{t}^{(S)}X\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{t}^{(H)}Y\partial_{t}^{(S)}Y\color[rgb]{0,0,0}-\color[rgb]{0,0,0}\partial_{t}^{(H)}Z\partial_{t}^{(S)}Z.$ (45)
Thus we obtain
$\dot{S}$ $=$ $-\partial_{t}^{(H)}X\partial_{t}^{(S)}X-\partial_{t}^{(H)}Y\partial_{t}^{(S)}Y-\partial_{t}^{(H)}Z\partial_{t}^{(S)}Z\$(46)
$+$ $\left(\partial_{t}^{(S)}X\right)^{2}+\left(\partial_{t}^{(S)}Y\right)^{2}+\left(\partial_{t}^{(S)}Z\right)^{2},\$
$\dot{H}_{1}$ $=$ $\partial_{t}^{(H)}H_{1}+\partial_{t}^{(S)}H_{1},\$(47)
$\partial_{t}^{(H)}H_{1}$ $=$ $0.$ (48)
Antisymmetry forces the bracket to vanish when an argument is duplicated; hence each $H_{k}$ is conserved along the Hamiltonian part of the dynamics.
$\partial_{t}^{(S)}H_{1}$ $=$ $\{H_{1},X,Y\}_{\mathrm{NB}}\{S,X,Y\}_{\mathrm{NB}}+\{H_{1},Y,Z\}_{\mathrm{NB}}\{S,Y,Z\}_{\mathrm{NB}}+\{H_{1},Z,X\}_{\mathrm{NB}}\{Z,X,S\}_{\mathrm{NB}}\$(49)
$=$ $\{H_{1},X,Y\}_{\mathrm{NB}}\partial_{t}^{(S)}Z+\{H_{1},Y,Z\}_{\mathrm{NB}}\partial_{t}^{(S)}X+\{H_{1},Z,X\}_{\mathrm{NB}}\partial_{t}^{(S)}Y,$
$\{H_{1},X,Y\}_{\mathrm{NB}}=-k_{5}hBX+2k_{3}AY,$ (50)
$\{H_{1},Y,Z\}_{\mathrm{NB}}=\frac{k_{2}X^{2}}{2}-k_{5}hBZ,$ (51)
$\{H_{1},Z,X\}_{\mathrm{NB}}=-k_{1}AY-\frac{k_{2}Y^{2}}{2}+2k_{3}AZ,$ (52)
$\dot{H}_{2}=\partial_{t}^{(H)}H_{2}+\partial_{t}^{(S)}H_{2},$ (53)
$\partial_{t}^{(H)}H_{2}=0,$ (54)
$\partial_{t}^{(S)}H_{2}=\partial_{t}^{(S)}Z=2k_{3}AX-k_{5}BZ.$ (55)
The sign of the Hamiltonian time evolution can be determined as follows. The Hamiltonian part of the time evolution is given by
$\frac{dF}{dt}\Big|_{H}=-\{F,H_{1},\dots,H_{n-1}\}_{\mathrm{NB}},$ (56)
where the Hamiltonian flow is defined as
$(v_{H})_{i}=\sum_{jk\dots l}\epsilon_{ijk\dots l}\frac{\partial H_{1}}{\partial x_{j}}\frac{\partial H_{2}}{\partial x_{k}}\dots\frac{\partial H_{n-1}}{\partial x_{l}}.$ (57)
The sign of the Hamiltonian contribution to the entropy rate is determined by
$\mathrm{sign}\left(\frac{dS}{dt}\Big|_{H}\right)=\mathrm{sign}\big(-\nabla S\cdot v_{H}\big),$ (58)
being positive when $v_{H}$ is aligned with $\nabla S$, negative when opposed, and zero when orthogonal.
In this paper the time evolution of the BZ reaction is numerically studied, which yields the following results depicted in Figures 1, 2, 4, 4, and 5. These figures illustrate how the BZ reaction oscillates, with the time evolution computed numerically using the Dormand–Prince method. The numerical stability was verified by tightening the relative and absolute tolerances and applying step-size control, confirming that no differences appeared in the trajectories or entropy diagnostics; the same verification was performed for the examples in the subsequent sections.
Figure 1 shows the temporal development of the concentrations, $X$, $Y$ and $Z$, as a function of time given in the horizontal axis. Figure 2 gives the time development of the entropy $S$.
Figure 4 gives the change of the Hamiltonian $H_{1},H_{2}$ and the entropy $S$ in time.
Figure 4 shows the temporal change of entropy $\frac{\partial S}{\partial t}$, $\frac{\partial^{(H)}S}{\partial t}$ and $\frac{\partial^{(S)}S}{\partial t}$ in detail.
As can be seen from Figure 4, the respective contributions from $\frac{\partial^{(H)}S}{\partial t}$ and $\frac{\partial^{(S)}S}{\partial t}$ almost cancel each other, and entropy is found to be almost unchanged, except for the sudden increase and decrease of the entropy, subjected to periodic delta-function-type positive and negative kicks. The positive and negative kicks arise alternately, where the period of positive kick $T_{+}$ and negative kick $T_{-}$ are both $T_{+}=T_{-}\approx 20$. It should also be noted that the contribution of the Hamiltonian part to the entropy transformation is always negative.
This can be seen also in the path diagram, Figure 5, where the orbits undergo an inverted kick at both ends of the elongated circle, and only then does entropy undergo an abrupt change. It can be seen from the figure that $H_{1}$ and $S$ are both zero at such a turning point.
Now, we can understand well why the usual non-equilibrium thermodynamics by Onsager can not be applied to the BZ reaction. The Onsager theory is formulated near the equilibrium point of the entropy. There, the entropy gradient describes the time evolution of the thermodynamic variable as an affinity force, and as time passes, entropy increases while its rate becomes zero. In systems far from equilibrium, dynamics are described by entropy gradients plus Nambu dynamics. In systems such as the BZ reaction, entropy does not necessarily increase, but rather oscillates between the Hamiltonians $H_{1}$ and $H_{2}$.
Figure 1: Time variation of concentration X, Y and Z. Horizontal axis represents time, vertical axis represents concentration.
Figure 2: Plot of entropy change over time. The horizontal axis represents time and the vertical axis represents entropy.
Figure 3: Plot of $H_{1}$, $H_{2}$ and $S$ as a function of time. The horizontal axis is time and the vertical axis is the values of $H_{1}$, $H_{2}$ and $S$.
Figure 4: Plot of entropy rate as a function of time. $\partial^{(H)}_{t}S$ and $\partial^{(S)}_{t}S$ represent the contribution of the Hamiltonian in the entropy rate and the dissipation due to entropy in the time variation of entropy, respectively, and $dS/dt$ is the sum of these contributions.
Refer to caption Refer to caption
Figure 5: 3D diagram of the limit cycle according to the time variation of the concentration X, Y and Z. The vertical axis represents the concentration of Z, the horizontal axis represents the concentration of X and the remaining axis represents the concentration of Y.
Note that oscillations do not occur in all parameter regions. In particular, when the parameter relation $hk_{5}\sim k_{1}A\sim k_{2}\sim k_{4}X\sim k_{5}B\ll k_{3}A$ holds, the BZ reaction exhibits oscillatory behavior. As discussed in Appendix C, under this condition the system can be approximated by a set of harmonic oscillator Hamiltonians, which provides a clear theoretical basis for the emergence of oscillations. Approximating the Hamiltonian and entropy in the case, we get
$H_{1}\approx 2k_{3}AYZ+\dots$ (59)
$H_{2}=Z,$
$S\approx\frac{1}{2}k_{3}AX^{2}+2k_{3}AXZ+\dots.$
This represents a system of harmonic oscillators as discussed in Appendix C. Thus, under the appropriate variable transformations, the system is approximately described as a system of harmonic oscillator Hamiltonians with no entropy. Thus, from the given Hamiltonian and entropy, we can analyze whether oscillatory phenomena occur [^5] When the condition $dS=dH_{1}\wedge d\Delta H_{2}+d\Delta H_{1}\wedge dH_{2}+d\Delta H_{1}\wedge d\Delta H_{2}$ is satisfied, the redefinition $H_{i}\rightarrow H_{i}+\Delta H_{i}$ leads to the elimination of the entropy $S$, leaving only the Hamiltonians to describe the system. This removal of $S$ is crucial in achieving a quasi-conserved system of Hamiltonians. In general, the interaction between the set of Hamiltonians and entropy plays a significant role in determining the system’s behavior..
3.2 Hindmarsh-Rose model
As a prototypical spiking system, we reconstruct the Hindmarsh–Rose (H–R) neuron model within NNET and discuss the time dependence of ($H_{i}$) and ($S$) across spike–burst cycles. The H-R model is a biological neuron model of the spike-burst behavior of the membrane potential which is a typical behavior of neurons in the brain[31][35]. Understanding spikes in the H–R model enables us to describe the fundamental mechanism of neural signal transmission within the framework of dynamical systems, and moreover, to place it in a unified discussion with phenomena such as the BZ reaction as a universal behavior of non-equilibrium systems. The H-R model is a three-dimensional dynamical system, which is given by the following equations:
$\dot{x}$ $=$ $y+\phi(x)-z+I,$ (60)
$\dot{y}$ $=$ $\psi(x)-y,$ (61)
$\dot{z}$ $=$ $r(s(x-x_{R})-z)$ (62)
where
$\phi(x)$ $=$ $-ax^{3}+bx^{2},$ (63)
$\psi(x)$ $=$ $c-dx^{2}$ (64)
where $x$ is the biological membrane potential and $x_{R}$ is the resting potential. $y$ is called the recovery variable as it takes into account the transport of ions across the membrane through the ion channel. The recovery variable $y$ represents slower ionic processes that counteract depolarization, enabling the membrane potential to return toward rest after a spike. $z$ is the bursting variable and $I$ is the current entering the neuron from outside and is used as a control parameter. $a,b,c,d,s$ are fixed parameters of the model and often take values such as $a=1,b=3,c=1,d=5,s=4$. In addition, $r$ represents the timescale related to neural adaptation and assumes a very small value on the order of $10^{-3}$. The time evolution has been numerically integrated by means of the Dormand–Prince method, as in the case of the BZ reaction.
From the same discussion as for the BZ reaction, we obtain
$\frac{\partial^{2}H_{1}}{\partial X^{2}}+\frac{\partial^{2}H_{1}}{\partial Y^{2}}$ $=$ $1+2dX,$ (65)
$-\frac{\partial^{2}H_{1}}{\partial Z\partial X}$ $=$ $0,$ (66)
$\frac{\partial^{2}H_{1}}{\partial Z\partial Y}$ $=$ $1+rs.$ (67)
From this,
$\frac{\partial H_{1}}{\partial X}$ $=$ $+f(X,Y),$ (68)
$\frac{\partial H_{1}}{\partial Y}$ $=$ $(1+rs)Z+g(X,Y),$ (69)
$\frac{\partial f(X,Y)}{\partial X}+\frac{\partial g(X,Y)}{\partial X}$ $=$ $1+2dX.$ (70)
By substituting these into the above formula,
$f(X,Y)$ $=$ $-dX^{2},\$(71)
$g(X,Y)$ $=$ $-Y.$ (72)
Therefore, $H_{1}$ is obtained as
$H_{1}=-\frac{d}{3}X^{3}-\frac{1}{2}Y^{2}+(1+rs)YZ.$ (73)
In the NNET, $H_{1}$, $H_{2}$, and $S$ corresponding to this model have the following forms:
$H_{1}$ $=$ $-\frac{d}{3}x^{3}-\frac{1}{2}y^{2}+(1+rs)yz,$ (74)
$H_{2}$ $=$ $z,$ (75)
$S$ $=$ $-\frac{a}{4}x^{4}+\frac{b}{3}x^{3}+Ix-\frac{y^{2}}{2}+cy-\frac{rz^{2}}{2}-rsx_{R}z+rsxz.$ (76)
Here we set $H_{2}=z$, since the bursting variable $z$ evolves on a much slower timescale ($r\ll 1$) and can thus be regarded as a pseudo-conserved quantity, analogous to the catalyst concentration in the BZ reaction. Also, we divide time evolution into Hamiltonian part and entropy part as follows:
$\partial_{t}^{(H)}x$ $=$ $-\{H_{1},H_{2},x\}_{\mathrm{NB}}=y-(1+rs)z,$ (77)
$\partial_{t}^{(H)}y$ $=$ $-\{H_{1},H_{2},y\}_{\mathrm{NB}}=-dx^{2},$ (78)
$\partial_{t}^{(H)}z$ $=$ $-\{H_{1},H_{2},z\}_{\mathrm{NB}}=0,$ (79)
$\partial_{t}^{(S)}x$ $=$ $I+bx^{2}-ax^{3}+rsz,$ (80)
$\partial_{t}^{(S)}y$ $=$ $c-y,$ (81)
$\partial_{t}^{(S)}z$ $=$ $r(s(x-x_{R})-z).$ (82)
Next, the time evolution of entropy is given by
$\dot{S}=\partial_{t}^{(H)}S+\partial_{t}^{(S)}S,$ (83)
where
$\partial^{(H)}_{t}S$ $=$ $-dx^{2}(c-y)-(I+bx^{2}-ax^{3}+rsz)(-y+(1+rs)z),$ (85)
$\partial^{(S)}_{t}S$ $=$ $(c-y)^{2}+r^{2}(s(x_{R}-x)+z)^{2}+(I+(b-ax)x^{2}+rsz)^{2}.$ (86)
The time evolution of Hamiltonian is given by
$\dot{H_{1}}$ $=$ $-r(1+rs)y(s(x-x_{R})-z)+dx^{2}(I+x^{2}(b-ax)+rsz)$ (87)
$+$ $(c-y)(y-(1+rs)z),$
$\dot{H_{2}}$ $=$ $r(s(x_{R}-x)+z).$ (88)
Figure 6: Time variation of membrane potential $x$, recovery variable $y$ and bursting variable $z$. Horizontal axis represents time, vertical axis represents $x$, $y$, $z$.
Figure 7: 3D diagram of the limit cycle according to the time variation of $x$, $y$, and $z$. The vertical axis represents the bursting variable $z$, the horizontal axis represents the membrane potential $x$ and the remaining axis represents the recovery variable $y$.
Figure 8: Plot of $H_{1}$, $H_{2}$, and $S$ as a function of time. The horizontal axis is time and the vertical axis is the values of $H_{1}$, $H_{2}$, and $S$.
Figure 9: Plot of entropy rate as a function of time. $\partial^{(H)}_{t}S$ and $\partial^{(S)}_{t}S$ represent the contribution of the Hamiltonian in the entropy rate and the dissipation due to entropy in the time variation of entropy, respectively, and $dS/dt$ is the sum of these contributions.
The spike-like excitatory activity of the membrane potential is shown in Figure 6 and Figure 7. The trajectory shows a limit cycle.
Figure 9 shows the time evolution of $H_{1}$, $H_{2}$, and $S$. It can be seen that $H_{1}$ and $S$ take alternating peaks in a spike-like pattern, while $H_{2}$ increases in a staircase-like pattern.
Figure 9 decomposes the time evolution of the rate of entropy into a contribution from the Hamiltonian and a contribution from entropy, with the Hamiltonian contribution of entropy having a periodic contribution to entropy.
We consider the conditions under which the time derivatives of $H_{1},\ H_{2},$ and $S$ are zero.
$\dot{H}_{2}$ $=r\{\,s(x_{R}-x)+z\,\}=0\ \Longleftrightarrow\ z=s(x-x_{R}),$ (89)
$\dot{S}$ $=0\ \Longleftrightarrow\ y=c,\quad z=s(x-x_{R}),\quad I+(b-ax)x^{2}+rs\,z=0,$ (90)
$\dot{H}_{1}$ $=0\ \Longleftrightarrow\ y=c,\quad z=s(x-x_{R}),\quad I+(b-ax)x^{2}+rs\,z=0.$ (91)
Eq. (89) is the $z$-nullcline (balance line of the bursting variable).
Intersections of Eq. (89)–Eq. (91) satisfy $\dot{H}_{1}=\dot{H}_{2}=\dot{S}=0$ and mark phase–flip points on the limit cycle (turning points or extrema of the spike velocity). Moreover, when the system is near this intersection, $H_{1}$ behaves as a pseudo-conserved quantity.
About $H_{2}$ increasing, $H_{2}=z,\ \dot{H}_{2}=r\{\,s(x_{R}-x)+z\,\},$ with $r\ll 1$, so that $z$ acts as a slow variable. The sign of $\dot{H}_{2}$ depends on the balance between $z$ and $s(x-x_{R})$; when $z>s(x-x_{R})$, which typically occurs during spiking with $x>x_{R}$, we obtain $\dot{H}_{2}>0$. This mechanism leads to the staircase-like increase of $H_{2}$, representing neuronal adaptation as a slow negative feedback.
3.3 Lorenz System
As paradigmatic chaotic flows, we reconstruct the Lorenz and Chen systems within NNET and discuss the time dependence of ($H_{i}$) and ($S$) along trajectories and across regime transitions. In this subsection, we analyze the Lorenz system[32, 36, 37], a representative dynamical system exhibiting chaotic behavior. The Lorenz system is widely known as a simplified model of convection and fluid dynamics.
The Lorenz system can be described as a toy model consisting of the following set of three differential equations:
$\dot{x}$ $=$ $\sigma(y-x),$ (92)
$\dot{y}$ $=$ $x(\rho-z)-y,$ (93)
$\dot{z}$ $=$ $xy-\beta z,$ (94)
where we set the parameters as $\sigma=10$, $\beta=\frac{8}{3}$ with the time evolution computed numerically using the Dormand–Prince method.
For $\rho<24.74$, a nonzero equilibrium exists and is stable, with trajectories eventually converging to it after damped oscillations. At $\rho\approx 24.74$, a bifurcation occurs, leading to the emergence of a stable periodic solution. When $24.74<\rho$, the periodic solution becomes unstable, and the system transitions to chaos.
Figure 10: Trajectory of the Lorenz system with initial condition $x(0)=1.0$, $y(0)=1.0$, $z(0)=1.0$, and $dt_{init}=0.005$. Displayed side by side for increasing values of $\rho$ ($\rho=13.0,\ 14.0,\ 15.0,\ 24.74,\ 28.0$): the leftmost panel shows convergence to the origin or weak oscillations, the middle panels display periodic orbits (either confined to one lobe or alternating between lobes), and the rightmost panel exhibits fully developed chaos.
Figure 10 shows the typical attractor of the Lorenz system in three-dimensional phase space. The two circular lobes correspond to trajectories revolving around the two nonzero equilibria. Chaos arises when the switching between these lobes becomes aperiodic, which can be quantified in the NNET framework: on the Poincaré section, the sequence $\{H_{1}(t_{k}),S(t_{k})\}$ evolves from clustered to band-like and eventually to diffuse distributions.
Within the framework of NNET, we introduce the following expressions for the Hamiltonians $H_{1}$, $H_{2}$, and the entropy SS [^6] Axenides and Floratos (A&F)[38] study dissipative dynamics by adopting a Helmholtz decomposition of the vector field (rotational part = non-dissipative, gradient part = dissipative), which is formally compatible with the type of splitting used in our discussion. Our NNET framework, however, is formulated as a non-equilibrium thermodynamics: the irreversible contribution is required to take the thermodynamic form $g^{ij}\nabla_{j}S$ with a non-negative dissipation metric, so that the structure of $\dot{S}$ explicitly separates entropy exchange from non-negative entropy production. In contrast, the gradient term in A&F is introduced primarily to represent phase-space volume contraction and is not tied to such thermodynamic constraints. Finally, the reversible/irreversible separation itself is generally non-unique; in NNET this is reflected, for example, in the reparameterization freedom of $H_{1},H_{2}$ (e.g., an $SL(2,\mathbb{R})$-type freedom). More concretely, this redundancy allows one to reshuffle contributions between the Nambu and gradient parts by a compensating transformation of the form $0=\{x^{i},H_{1},Q\}_{\mathrm{NB}}+\partial_{i}(\delta S)$ (for a suitable pair $(Q,\delta S)$), which leaves $\dot{x}^{i}$ unchanged while modifying the split. Accordingly, our representation $(H_{1},H_{2},S)$ is written so as to retain this redundancy.:
$H_{1}$ $=$ $\frac{1}{2}(-\rho x^{2}+\sigma y^{2}+\sigma z^{2}),$ (95)
$H_{2}$ $=$ $z-\frac{1}{2\sigma}x^{2},$ (96)
$S$ $=$ $-\frac{1}{2}(\sigma x^{2}+y^{2}+\beta z^{2}).$ (97)
Figure 11: Time evolution of the Hamiltonian $H_{1}$ for the Lorenz system over the time interval $0\leq t\leq 60$ and $dt_{init}=0.005$ with initial condition $x(0)=1.0$, $y(0)=1.0$, and $z(0)=1.0$.
Figure 12: Time evolution of the Hamiltonian $S$ for the Lorenz system over the time interval $0\leq t\leq 60$ and $dt_{init}=0.005$ with initial condition $x(0)=1.0$, $y(0)=1.0$, and $z(0)=1.0$.
We next apply the NNET section analysis to the Lorenz system. Figures 11 and 12 display the time evolution of $H_{1}$ and $S$ over the interval $0\leq t\leq 60$ with initial condition $x(0)=y(0)=z(0)=1.0$ and $dt_{init}=0.005$. To extract the underlying structure, we sample the trajectory at times $t_{k}$ when it intersects the section $z=z^{\ast}:=\rho-1$, corresponding to the $z$-coordinate of the nonzero equilibria, and record the values $H_{1}(t_{k})$ and $S(t_{k})$.
The sampled sequences reveal a clear progression: in the stationary regime the points remain tightly clustered, in the periodic regime they form band-like structures reflecting alternation between the two lobes, and in the chaotic regime they spread broadly across the section. Thus, the two circular lobes in the phase space are directly connected to these rearrangements in the $(H_{1},S)$ plane.
Moreover, near bifurcation points the fluctuation range of $S(t_{k})$ grows significantly, and the clustering of points becomes asymmetric due to biased residence times in each lobe. This behavior highlights how the onset of chaos is accompanied by measurable changes in the distributions of $H_{1}$ and $S$. Consequently, even from section samples alone, NNET enables a systematic classification of stationary, periodic, and chaotic states, providing a unified dynamical interpretation.
3.4 Chen System
In this subsection, we analyze the Chen system[33], which exhibits behavior similar to the Lorenz system[32], a representative dynamical system with chaotic behavior. Unlike the Lorenz system, where the phase-space contraction rate is constant and bifurcations produce abrupt changes, the Chen system exhibits a parameter-dependent contraction rate, leading to continuous deformations of the phase-space structure. This manifests in the NNET section analysis as gradual distortions and asymmetries in the $(H_{1},S)$ distribution, rather than sharp transitions.
Like the Lorenz system, the Chen system can be described as a toy model consisting of the following set of three differential equations:
$\dot{x}$ $=$ $a(y-x),$ (98)
$\dot{y}$ $=$ $(c-a)x+cy-xz,$ (99)
$\dot{z}$ $=$ $xy-bz,$ (100)
where we set the parameters as $a=35$, $b=3$ with the time evolution computed numerically using the Dormand–Prince method. While the Lorenz system exhibits a constant rate of phase-space volume contraction (uniformly dissipative), the Chen system’s contraction rate varies with the parameter $c$, resulting in a continuous change.
For $17.5<c<20.08$, a nonzero equilibrium exists and is stable, with trajectories eventually converging to it after damped oscillations. At $c\approx 20.08$, a bifurcation occurs, leading to the emergence of a stable periodic solution. When $20.08<c<38$, the periodic solution becomes unstable, and the system transitions to chaos.
Figure 13: Trajectory of the Chen system with initial condition $x(0)=1.0$, $y(0)=1.0$, $z(0)=1.0$, and $dt_{init}=0.005$. Placed side by side for c=18.0, 19.0c=18.0,\ 19.0 (damped oscillations converging to equilibrium), $c\approx 20.08$ (emergence of a periodic orbit), and $c=28.0$ (chaos).
Within the framework of NNET, we introduce the following expressions for the Hamiltonians $H_{1}$, $H_{2}$, and the entropy $S$:
$H_{1}$ $=$ $\frac{1}{2}(2a-c)(y^{2}+z^{2}),$ (101)
$H_{2}$ $=$ $-z+\frac{1}{2(2a-c)}x^{2},$ (102)
$S$ $=$ $-\frac{1}{2}ax^{2}+\frac{1}{2}cy^{2}-\frac{1}{2}bz^{2}+(c-a)xy$ (103)
where 2ac>02a-c>0.
Figure 14: Time evolution of the Hamiltonian $H_{1}$ for the Chen system over the time interval $0\leq t\leq 60$ and $dt_{init}=0.005$ with initial condition $x(0)=1.0$, $y(0)=1.0$, and $z(0)=1.0$.
Figure 15: Time evolution of the Hamiltonian $S$ for the Chen system over the time interval $0\leq t\leq 60$ and $dt_{init}=0.005$ with initial condition $x(0)=1.0$, $y(0)=1.0$, and $z(0)=1.0$.
Figure 13 shows the typical attractor of the Lorenz system in three-dimensional phase space. As in the case of the Lorenz system, in Figures 14, 15 we evaluate the intersection time series $H_{1}(t_{k})$ and $S(t_{k})$ on the section $z=z^{\ast}:=2c-a$, corresponding to the components of the zero equilibrium. As $c$ increases, the point sequence transitions from unimodal dispersion to multimodal and broad distribution, suggesting the progression from the destabilization of periodic solutions to the chaotic phase. In particular, the band expansion of $H_{1}$ is consistent with the characteristic driving enhancement of the Chen system (the increase of 2ca2c-a).
It is worth noting that in the Lorenz system, dissipation is constant, so the time evolution of $S$ is highly sensitive to changes in the threshold value of $\rho$, and the $(H_{1},S)$ patterns change dramatically at each bifurcation. In contrast, in the Chen system the dissipation rate depends explicitly on $c$, and thus the variations in $(H_{1},S)$ evolve more continuously, appearing as asymmetry in orbital residence times and clustering on the attractor.
4 Summary and Discussion
We have shown that Nambu Non-equilibrium Thermodynamics (NNET) gives a practical and unified description of far-from-equilibrium dynamics across four paradigmatic systems: the BZ reaction, the Hindmarsh–Rose neuron, and the Lorenz and Chen chaotic flows. In all cases the decomposition into multi-Hamiltonian (incompressible) flow and an entropic (gradient) flow cleanly separates structure-forming dynamics from dissipation.
For the BZ reaction, the dynamics require multiple Hamiltonians together with an explicit entropic term. Along a cycle, entropic and Hamiltonian contributions almost cancel, concentrating entropy changes into short “kicks”; this clarifies why Onsager’s near-equilibrium theory fails here and how oscillations persist despite dissipation. For the Hindmarsh–Rose model, the zero-set conditions for $\dot{H}_{1},\dot{H}_{2},\dot{S}$ identify phase-flip points on the limit cycle. Near these intersections, $H_{1}$ behaves as a pseudo-conserved quantity, which accounts for the timing of spikes, while the slow dynamics of $H_{2}=z$ with $\dot{H}_{2}=r{s(x_{R}-x)+z}$ explain its staircase-like increase.
For Lorenz (constant phase-space contraction) and Chen (contraction rate depending on $c$), an NNET section analysis of $\{H_{1}(t_{k}),S(t_{k})\}$ reveals a common progression—clustered $\rightarrow$ band-like $\rightarrow$ diffuse—tracking stationary, periodic, and chaotic regimes; in Chen this progression deforms continuously with $c$, producing residence-time asymmetries, whereas in Lorenz the patterns change sharply at bifurcations.
Although we focused on spatially uniform ODEs, the construction readily extends to spatially extended systems. This suggests new analyses of entropy–Hamiltonian interplay in pattern formation (e.g., Turing[16] and reaction–diffusion[39, 40]) and other oscillatory or chaotic media (e.g., Bray–Liebhafsky (BL) model[41], lasers[42], astrophysics[43], earthquakes[44], engineering applications[45, 46], and psychology[47]). We hope that these results will eventually contribute to a unified understanding of spike transmission in neuroscience or to a new classification principle for chemical reaction systems; these remain important subjects for future work.
Acknowledgments
We would like to thank Toshio Fukumi about non-linear response theory. We are indebted to Shiro Komata for reading this paper and giving useful comments.
Appendix Appendix A NNET in summary (for the present 3D applications)
We consider an $n$-dimensional state vector $\bm{x}=(x^{1},\dots,x^{n})$ on a regular region of the state space. Following the axiomatic formulation of NNET[25], the time evolution is decomposed into a reversible part and an irreversible part,
$\dot{x}^{i}=v_{\rm rev}^{i}(\bm{x})+v_{\rm irr}^{i}(\bm{x}).$ (104)
The reversible part is generated by the Nambu bracket with multiple Hamiltonians, while the irreversible part is a gradient flow driven by an entropy-like potential $S$ with a positive definite coefficient (metric) $g^{ij}$.
In the 3D systems studied in this paper ($n=3$), the Nambu bracket is the Jacobian
$\{F,G,H\}:=\frac{\partial(F,G,H)}{\partial(x,y,z)}=\nabla F\cdot(\nabla G\times\nabla H).$ (105)
The NNET evolution law used throughout the main text reads
$\dot{\bm{x}}=\{\bm{x},H_{1},H_{2}\}+\sum_{j=1}^{3}g^{ij}\,\nabla_{j}S,$ (106)
(or its component form with $g^{ij}$). Here “reversible” is used in the thermodynamic-structure sense: the Nambu part is a volume-preserving (incompressible) flow and conserves the Hamiltonians in the absence of dissipation; it is not meant to imply strict time-reversal invariance under all physical constraints (e.g., positivity constraints of variables).
The entropy rate is decomposed as
$\dot{S}=\{S,H_{1},H_{2}\}+\sum_{i,j=1}^{3}(\nabla_{i}S)^{\mathsf{T}}g^{ij}(\nabla_{j}S).$ (107)
The second term is non-negative for positive definite $g^{ij}$ and corresponds to entropy production. The first term is not sign-definite and represents entropy exchange/flux; hence $\dot{S}$ can be negative in open or far-from-equilibrium situations.
The decomposition in Eq. (106) is not unique in general. (i) The Hamiltonians admit reparameterizations $(H_{1},H_{2})\mapsto(\Phi(H_{1},H_{2}),\Psi(H_{1},H_{2}))$. (ii) One may shift contributions between the Nambu and gradient parts unless additional conditions are imposed.
In this work we fix the split by the following practical rule: (1) choose $(H_{1},H_{2})$ so that the Nambu part reproduces the divergence-free component of the given vector field, (2) choose $S$ so that the remaining component is captured by a gradient flow with a positive definite $g^{ij}$, (3) when multiple choices exist, select the simplest polynomial (or lowest-order) $(H_{1},H_{2},S)$ consistent with the model parameters.
Appendix Appendix B Analysis of cycles and spikes
Two characteristic behaviors of systems far from equilibrium are cyclic trajectories and the occurrence of spikes. When conserved quantities exist, the trajectories are expected to form cycles, as they must continually evolve to satisfy the conservation laws. Even in the absence of exact conserved quantities, cycles can persist over long timescales if there exist pseudo-conserved quantities that decrease only asymptotically. The triangular reaction[25] has already been presented as an example of such a system with a conserved quantity; here, however, we also consider the following simple toy model.
$H_{1}=a-x^{3}z,$ (108)
$H_{2}=b+z,$ (109)
$S=x^{2}yz.$ (110)
The equations of motion are as follows
$\dot{x}=2xyz,$ (111)
$\dot{y}=-2x^{2}z,$ (112)
$\dot{z}=x^{2}y.$ (113)
This preserves
$O=\frac{1}{2}(x^{2}+y^{2})$ (114)
in time evolution. The path draws a limit cycle as in Figures 17 and 17, with a sharp spike for the $x$ and $z$ components resulting from its nonlinearity.
Figure 16: 3D diagram of the limit cycle according to the time variation of $x$, $y$, and $z$.
Figure 17: Plot of $x$, $y$ and $z$ as a function of time.
Another toy model is
$H_{1}=H_{1}(x,y,z),$ (115)
$H_{2}=(x^{2}-y^{2})f(z),$ (116)
$S=xy,$ (117)
where $H_{1}(x,y,z)$ and $f(z)$ are arbitrary functions. The equations of motion are
$\dot{x}=y-2yf(z)\frac{\partial H_{1}}{\partial z}+(y^{2}-x^{2})\frac{df(z)}{dz}\frac{\partial H_{1}}{\partial y},$ (118)
$\dot{y}=x-2xf(z)\frac{\partial H_{1}}{\partial z}-(y^{2}-x^{2})\frac{df(z)}{dz}\frac{\partial H_{1}}{\partial x},$ (119)
$\dot{z}=2xf(z)\frac{\partial H_{1}}{\partial y}+2yf(z)\frac{\partial H_{1}}{\partial x}.$ (120)
In this model, $H_{2}$ itself is the conserved quantity of time evolution.
As a concrete example, we take
$H_{1}=\frac{1}{2}(x^{2}+y^{2}+z^{2}),$ (121)
$f(z)=z^{2}.$ (122)
The equations of motion are
$\dot{x}=y-2x^{2}yz+2y^{3}z-2yz^{3},$ (123)
$\dot{y}=x+2x^{3}z-2xy^{2}z-2xz^{3},$ (124)
$\dot{z}=4xyz^{2}.$ (125)
The path draws a limit cycle as in Figures 19 and 19, with a sharp spike for the $y$ and $z$ components resulting from its nonlinearity.
Figure 18: 3D diagram of the limit cycle according to the time variation of $x$, $y$, and $z$.
Figure 19: Plot of $x$, $y$, and $z$ as a function of time.
Appendix Appendix C Harmonic oscillators
At the end of Subsection 3.1, it was noted that, for a particular choice of parameters, the BZ reaction can be represented by a system of harmonic oscillators. In this Appendix, we examine this simplified model, which is characterized by the following Hamiltonians and entropy:
$H_{1}$ $=aYZ,$ (126)
$H_{2}$ $=Z,$ (127)
$S$ $=\frac{1}{2}bX^{2}+cXZ.$ (128)
The evolution of the system follows the Nambu non-equilibrium thermodynamic equation:
$\dot{X}^{i}$ $=-\{X^{i},H_{1},H_{2}\}+\frac{\partial S}{\partial X^{i}}.$ (129)
For the variables $X$, $Y$, and $Z$, this gives the following equations of motion:
$\dot{X}$ $=-aZ+bX+cZ=(c-a)Z+bX,$ (130)
$\dot{Y}$ $=0,$ (131)
$\dot{Z}$ $=cX.$ (132)
We can express this system in matrix form as:
$\begin{pmatrix}\dot{X}\\ \dot{Z}\end{pmatrix}$ $=\begin{pmatrix}b&c-a\\ c&0\end{pmatrix}\begin{pmatrix}X\\ Z\end{pmatrix}$. (133)
Differentiating once more gives:
$\begin{pmatrix}\ddot{X}\\ \ddot{Z}\end{pmatrix}=\begin{pmatrix}b&c-a\\ c&0\end{pmatrix}^{2}\begin{pmatrix}X\\ Z\end{pmatrix}=\begin{pmatrix}b^{2}+c(c-a)&b(c-a)\\ bc&c(c-a)\end{pmatrix}\begin{pmatrix}X\\ Z\end{pmatrix}$. (134)
The eigenvalue equation of this matrix,
$\lambda^{2}-\bigl(b^{2}+2c(c-a)\bigr)\lambda+c^{2}(c-a)^{2}=0.$ (135)
Since $\lambda_{+}\lambda_{-}=\det(M^{2})=c^{2}(c-a)^{2}\geq 0$, both eigenvalues have the same sign. In particular, when $b^{2}+2c(c-a)<0$ (i.e. $c(c-a)<-b^{2}/2$), both eigenvalues are negative, and the corresponding eigenmode is oscillatory (elliptic). In other words, there exists a conserved quantity, $\tilde{H}_{1}$, expressed as a quadratic algebraic relation in $(X,Z)$.
Since $Y$ is known to be another conserved quantity, which may be denoted by $\tilde{H}_{2}$, the system is thereby endowed with two conserved quantities, $(\tilde{H}_{1},\tilde{H}_{2})$. With this new choice of conserved quantities, no dissipation appears, implying $\tilde{S}=0$, and the dynamics is governed purely by the Nambu formalism.
This simplified model illustrates that even when entropy $S$ is present, an alternative selection of Hamiltonians and entropy may reveal hidden conservation laws within dissipative non-equilibrium thermodynamics. This phenomenon, referred to in this paper as “quasi-periodicity” or the emergence of a “quasi-cycle,” highlights the intriguing possibility that behaviors far from equilibrium can be systematically analyzed within the framework of Nambu non-equilibrium thermodynamics.
Appendix Appendix D Turning points of spikes
In this section, we discuss the turning points of spikes in the toy model in Section 2.
From Eq. (123)–(125), we obtain the following equations,
$dt$ $=\frac{dx}{y-2x^{2}yz+2y^{3}z-2yz^{3}}$ (136)
$=\frac{dy}{x+2x^{3}z-2xy^{2}z-2xz^{3}}$ (137)
$=\frac{dz}{4xyz^{2}}.$ (138)
Now consider a cusp that changes abruptly only in the $x$ direction. Since the denominator is quadratic for $x$, the equation can be rearranged as follows.
$dt=\frac{dx}{(x-\alpha_{+})(x-\alpha_{-})}=dx\left(\frac{A_{+}}{x-\alpha_{+}}+\frac{A_{-}}{x-\alpha_{-}}\right),$ (139)
$\alpha_{\pm}\equiv\mp\sqrt{y^{2}-z^{2}+\frac{1}{2z}},$ (140)
$A_{\pm}=2\alpha_{\pm}.$ (141)
Since $dt$ is singular in terms of the cusp arising and only the contribution of the $A_{+}$ term works around $x=\alpha_{+}$, we can write the following equation.
$\frac{x(t)-\alpha_{+}}{x(t_{0})-\alpha_{+}}=e^{(t-t_{0})/A_{+}}=(t-t_{0})/A_{+}+O((t-t_{0})^{2}/A_{+}^{2})$ (142)
Therefore, the gap before and after the turning point can be estimated from the following equation:
$\dot{x}(t_{*}+\delta t)-\dot{x}(t_{*}-\delta t)$ $=1/A_{+}\left(x(t_{*}+\delta t)+x(t_{*}-\delta t)\right)$ (143)
$=\frac{1}{2\sqrt{y(t_{*})^{2}-z(t_{*})^{2}+\frac{1}{2z(t_{*})}}}\left(x(t_{*}+\delta t)+x(t_{*}-\delta t)\right),$ (144)
where $t_{*}$ is time at the turning point.